OpenStudy (anonymous):

1/f = 1/p + 1/q solve for p.

OpenStudy (anonymous):

So we have: \[\frac{1}{f}-\frac{1}{q}=\frac{1}{p} \]Multiplying by \(pf\): \[ p-f=\frac{pf}{q} \]Dividing by \(pf\): \[ \frac{p-f}{pf}=\frac{1}{q} \]And taking the reciprocal: \[ \frac{pf}{p-f}=q \]And we are done.

Parth (parthkohli):

\[{1 \over f} = {1 \over p }+{1 \over q} \]\[\implies {1 \over f} = {q + p \over pq}\]\[\implies {pq \over f} = { q + p} \]\[\implies pq = f(p + q)\]\[\implies pq = fp + fq\]\[\implies pq - fq= fp\]\[\implies {q(p - f) \over f} = p \]

Parth (parthkohli):

@LolWolf Unfortunately, all that was done for \(q\). :P

OpenStudy (anonymous):

Oh, I solved for the wrong variable, but you can switch them, equally XD

OpenStudy (anonymous):

Haha, yeah, @ParthKohli

OpenStudy (anonymous):

Oy, be careful @ParthKohli , you still have a remaining \(p\) in your final equation.

Parth (parthkohli):

Oh... what the heck

OpenStudy (anonymous):

@emb27 I solved for the wrong variable, but you can switch them, equally, they are the same thing.

Parth (parthkohli):

\[\implies {q(p - f) \over f} ={pf \over f}\]Equating numerators:\[\implies pq - pf = pf\]\[\implies pq = 2pf\]\[\implies p = {2pf \over q}\]

Parth (parthkohli):

Is that correct? I need to sleep!

OpenStudy (anonymous):

I don't think it is, because we can do: \[ \frac{1}{f}-\frac{1}{q}=\frac{1}{p} \](Won't solve for the wrong variable this time!) To equal: \[ \frac{f-q}{fq}=\frac{1}{p} \]So: \[ \frac{1}{\frac{f-q}{fq}}=\frac{1}{\frac{1}{p}}=p=\frac{fq}{f-q} \]And you still have a remaining \(p\) on the other side XD

OpenStudy (anonymous):

Honestly, I am so confused. Which is the correct answer to 1/f = 1/p + 1/q. Solve for p.

OpenStudy (anonymous):

The correct answer is: \[ p=-\frac{fq}{f-q} \]

OpenStudy (anonymous):

Is the answer a positive or negative, both are shown?

OpenStudy (anonymous):

Oh, dang, I forgot to put a negative sign in my answer? What is wrong with me, to-day? Lack of sleep, I guess... yes, it's negative. I've messed up on the explanation, but the final case is correct.